Find the x- and y-intercepts and graph the equation plotting the intercepts
5x-6y=-45

Answer by rfer(16322)

5x-6y=-45
5x-6*0=-45
x=-9
(-9,0)
--------------------
5*0-6y=-45
-6y=-45
y=7.5
(0,7.5)